OpenStudy (anonymous):

Show would you go about solving this? Solve for y: x=(y-7/2)^1/5+3

OpenStudy (anonymous):

One way to think about solving this is to "simplify" the right hand side until it is just y, then what's on the left will be equal to y. So, first let's get rid of that 3.\[x-3=(y-\frac{7}{2})^\frac{1}{5}\] Now, what's the next thing to eliminate? Well, we could raise everything to the 5th power and eliminate that exponent.\[(x-3)^5=y-\frac{7}{2}\] Now, let's add 7/2, and we're done!\[y=(x-3)^5+\frac{7}{2}\]

OpenStudy (anonymous):

Thanks!